# PFGuard: A Generative Framework with Privacy and Fairness Safeguards

### v1

@@ Appendix F Related Work
Continuing from Sec. 6, we present more related work.
@@
Most privacy-preserving data generation techniques focus on satisfying differential privacy (DP) (Dwork et al., 2014). The majority of these techniques use Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) with DP training techniques, although privatizing other generative models have recently been proposed as well (Takagi et al., 2021; Cao et al., 2021; Harder et al., 2021; Liew et al., 2021; Chen et al., 2022; Vinaroz et al., 2022; Yang et al., 2023; Ghalebikesabi et al., 2023). One approach is based on DP-SGD (Abadi et al., 2016), which is a DP-enabled standard stochastic gradient descent algorithm to train ML models (Xie et al., 2018; Zhang et al., 2018; Torkzadehmahani et al., 2019; Bie et al., 2023). Another approach is based on the Private Aggregation of Teacher Ensembles (PATE) framework (Papernot et al., 2016; 2018), which trains multiple models on private data, and updates the generator with differentially private aggregation of multiple model outcomes (Jordon et al., 2018; Long et al., 2021; Wang et al., 2021a). GS-WGAN (Chen et al., 2020) is a state-of-the-art GAN-based technique that combines DP-SGD and PATE where multiple models are trained as in PATE, but their outcomes are processed with the Gaussian mechanism to update the generator as in DP-SGD. In comparison, PFGuard complements private GANs by also improving the fairness of data generation.
@@
The goal of model fairness is to avoid discriminating against certain demographics (Barocas et al., 2017; Feldman et al., 2015; Hardt et al., 2016), and fair data generation solves this problem by generating synthetic data to remove data bias. The main approaches of fair data generation are as follows: 1) modifying training objectives to balance model training (Xu et al., 2018; Sattigeri et al., 2019; Yu et al., 2020; Choi et al., 2020; Teo et al., 2023) and 2) modifying latent distributions of the input noise to obtain fairer outputs (Tan et al., 2020; Humayun et al., 2021). In comparison, PFGuard modifies sampling procedures to balance model training while preserving original training objectives and makes the key contribution of satisfying both privacy and fairness. There is another recent line of work using generated data together with original training data for model fairness (Roh et al., 2023; Zietlow et al., 2022), but they focus on classification tasks and assume to use given generative models.
@@
Recent studies have shown that achieving DP can hurt model fairness in classification tasks (Bagdasaryan et al., 2019; Farrand et al., 2020; Xu et al., 2020; Esipova et al., 2022), decision-making processes (Pujol et al., 2020), and even in generation tasks (Ganev et al., 2022). In addition, there is another notable line of work to investigate privacy-fairness-utility tradeoff, showing that achieving both privacy and fairness will necessarily sacrifice utility (Cummings et al., 2019; Agarwal, 2021; Sanyal et al., 2022). In comparison, our study uncovers the counteractive nature of privacy and fairness, which can result in compromised privacy and fairness, and ultimately compromised utility. To effectively achieve both privacy and fairness in model training, various techniques have been developed for classification tasks (Jagielski et al., 2019; Xu et al., 2019; 2020; Tran et al., 2022; Esipova et al., 2022; Lowy et al., 2023). In comparison, PFGuard focuses on data generation tasks, specifically tailoring its fair training phase to generative modeling objectives, which is learning the underlying training data distributions to generate synthetic data.

### v2

@@ Appendix F Related Work
Continuing from Sec. 6, we provide more discussion on related work.
@@
Most privacy-preserving data generation techniques focus on satisfying differential privacy (DP) (Dwork et al., 2014). The majority of these techniques focus on privatizing Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), while recent studies have explored other generative models as well (Takagi et al., 2021; Cao et al., 2021; Harder et al., 2021; Liew et al., 2021; Chen et al., 2022; Vinaroz et al., 2022; Yang et al., 2023; Ghalebikesabi et al., 2023). One privatization approach is based on DP-SGD (Abadi et al., 2016), which is a DP version of the standard stochastic gradient descent algorithm to train ML models (Xie et al., 2018; Zhang et al., 2018; Torkzadehmahani et al., 2019; Bie et al., 2023). Another privatization approach is based on the Private Aggregation of Teacher Ensembles (PATE) framework (Papernot et al., 2016; 2018), which trains multiple teacher models on private data, and updates the generator with differentially private aggregation of teacher model outcomes (Jordon et al., 2018; Long et al., 2021; Wang et al., 2021a). GS-WGAN (Chen et al., 2020) leverages both DP-SGD and PATE, where multiple teacher models are trained as in PATE, while their outcomes are processed with gradient cliiping as in DP-SGD. We design PFGuard to complement these private GANs by also achieving fairness in data generation.
@@
The goal of model fairness is to avoid discriminating against certain demographics (Barocas et al., 2017; Feldman et al., 2015; Hardt et al., 2016), and fair data generation solves this problem by generating synthetic data to remove data bias. The main approaches of fair data generation are as follows: 1) modifying training objectives to balance model training (Xu et al., 2018; Sattigeri et al., 2019; Yu et al., 2020; Choi et al., 2020; Teo et al., 2023) and 2) modifying latent distributions of the input noise to obtain fairer outputs (Tan et al., 2020; Humayun et al., 2021). In comparison, PFGuard 1) modifies sampling procedures to balance model training while preserving original training objectives and 2) makes the key contribution of satisfying both privacy and fairness. There is another recent line of work using generated data together with original training data for model fairness (Roh et al., 2023; Zietlow et al., 2022), but they focus on classification tasks and assume to use given generative models.
@@
Recent studies have shown that achieving DP can hurt model fairness in classification tasks (Bagdasaryan et al., 2019; Farrand et al., 2020; Xu et al., 2020; Esipova et al., 2022), decision-making processes (Pujol et al., 2020), and generation tasks (Cheng et al., 2021; Ganev et al., 2022; Bullwinkel et al., 2022; Rosenblatt et al., 2024). In addition, many studies have investigated the privacy-fairness-utility tradeoff, showing that achieving both privacy and fairness will necessarily sacrifice utility (Cummings et al., 2019; Agarwal, 2021; Sanyal et al., 2022). In comparison, our study uncovers the counteractive nature of privacy and fairness – achieving DP can compromise model fairness and achieving model fairness can compromise DP.
Private and Fair Classification
To effectively achieve both privacy and fairness in model training, various techniques have been developed for classification tasks (Jagielski et al., 2019; Xu et al., 2019; 2020; Tran et al., 2022; Esipova et al., 2022; Kulynych et al., 2022; Yaghini et al., 2023; Lowy et al., 2023). In comparison, PFGuard focuses on data generation tasks, which aim to learn the underlying training data distributions to generate synthetic data, and specifically tailors its fair training phase to generative modeling objectives. Although having the fundamental different problem settings, we make comparisons with existing classification techniques that also leverage 1) importance sampling or 2) PTEL, which are the two key components of PFGuard. (Kulynych et al., 2022) also extends importance sampling to private settings and evaluate fairness in downstream classification. In contrast, PFGuard evaluates fairness in both synthetic data generation and downstream classification. (Yaghini et al., 2023) and (Tran et al., 2022) use PTEL mechanisms, but rely on public datasets to train student classifiers. In contrast, PFGuard eliminates the need for public datasets by making PTEL queries with generated samples from the student generator. We finally note that  (Lowy et al., 2023) introduces the first DP fair learning method with convergence guarantees for empirical risk minimization. In contrast, PFGuard provides convergence guarantees for fair generative modeling.